Find the complete integral of (x+y)(p+q)^2+(x-y)(p-q)^2=1


1
Expert's answer
2020-09-21T09:59:43-0400

To solve differential equation "(x+y)(p+q)^2+(x-y)(p-q)^2 = 1"


Let us put "x+y = X^2" and "x-y = Y^2"

then

"p = \frac{\partial z}{\partial x} = \frac{\partial z}{\partial X}\frac{\partial X}{\partial x}+ \frac{\partial z}{\partial Y}\frac{\partial Y}{\partial x} = \frac{1}{2X}\frac{\partial z}{\partial X}+ \frac{1}{2Y}\frac{\partial z}{\partial Y}"


"q = \frac{\partial z}{\partial y} = \frac{\partial z}{\partial X}\frac{\partial X}{\partial y}+ \frac{\partial z}{\partial Y}\frac{\partial Y}{\partial y} = \frac{1}{2X}\frac{\partial z}{\partial X}- \frac{1}{2Y}\frac{\partial z}{\partial Y}"


Then

"p+q = \frac{1}{X}\frac{\partial z}{\partial X}"

and

"p-q = \frac{1}{Y}\frac{\partial z}{\partial Y}"


Now substitute it in the equation, then we get

"(\frac{\partial z}{\partial X } )^2 + (\frac{\partial z}{\partial Y } )^2 = 1"


Complete integral of the equation is

"z = aX+bY+c" where "a^2 + b^2 = 1"

Put value of X and Y, we get


"z = a\sqrt{(x+y)} + \sqrt{1-a^2}\sqrt{(x-y)} + c"


where a and c are arbitrary constants.